OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality y greater than or equal to 2x + 6 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

I believe it is the solid line without the origin....but someone else please correct me.

OpenStudy (anonymous):

i think it's C:)

OpenStudy (anonymous):

im with u @adam.bloedel lol

OpenStudy (johnweldon1993):

@LizzyLove<3 definitely not C nor A...since it is greater than or EQUAL to...it makes it a solid line...just not 100% sure on the shading lol

jimthompson5910 (jim_thompson5910):

To graph y >= 2x + 6, you graph the solid line y = 2x+6 Then you shade above the solid line

OpenStudy (johnweldon1993):

nvm you're right @adam.bloedel

jimthompson5910 (jim_thompson5910):

y >= 2x + 6 0 >= 2(0) + 6 0 >= 0 + 6 0 >= 6 which is FALSE, so (0,0) is NOT in the shaded region

OpenStudy (anonymous):

ohh sry!!! i tots messed tht up!!! i meant B or the second choice sry @johnweldon1993

OpenStudy (johnweldon1993):

lol no problem :) just glad you know the mistake :) and THANK YOU for that clarification @jim_thompson5910

OpenStudy (anonymous):

yupp:D @johnweldon1993